OpenStudy (anonymous):

are the given sides acute, right or obtuse: 4, 8, 6

OpenStudy (solomonzelman):

4+3+2=9 (parts) (4/9) x180=80 (3/9) x180=60 (2/9) x180=40 So the smallest angle=40 So your angles are 40, 60, 80

OpenStudy (solomonzelman):

give me your option about the classification.